OpenStudy (anonymous):

35 is what percent of 70? is the answer 50?

OpenStudy (ganpat):

70 * x% = 35 70 * (x/100) = 35 x/100 = 35/70 x = 1/2 * 100 x = 50 %...

OpenStudy (anonymous):

thank yu